OpenStudy (anonymous):

A water tank in the form of an inverted frustum of a cone has an altitude of 8 ft., and upper and lower radii of 6 ft. and 4 ft., respectively. Find the volume of the water tank and the wetted part of the tank if the depth of the water is 5 ft.

OpenStudy (anonymous):



OpenStudy (anonymous):

\(\Huge V=\frac{\pi h} {3}( R^2+Rr+r^2)\) such that \(V\) is the volume \(R\) the big raduis \(r\) the small raduis

OpenStudy (anonymous):

* and \(h\) is the hight \(\Huge V=\frac{8\pi}{3}(6^2+6×4+4^2)\)

OpenStudy (anonymous):

\(\Huge V= \frac{608\pi}{3}\)

OpenStudy (anonymous):

i hope that u get it , let me know if not :D